# A Large Recurrent Action Model: xLSTM Enables Fast Inference for Robotics Tasks

### v2

@@ 4.2 Scaling comparison
To conduct our main comparisons, we train our four backbone variants on the full training task mixture of 432 tasks. For each architecture backbone, we report performance scores for four model sizes: 16M, 48M, 108M, and 206M parameters. We train all models for 200K updates with a batch size of 128 and context length of 50 timesteps. All domains are represented with approximately equal proportion, resulting in 33K updates per domain. Additional implementation details and hyperparameters for every backbone variant and model size are available in Appendix C.
@@
Evaluation performance. During training, we evaluate our agents after every 50K step in all 432 training environments. In Figure 2b, we report the resulting normalized performances averaged across all six domains. The recurrent backbones outperform the Transformer one across model sizes. While xLSTM and Mamba performs similarly at smaller scales, xLSTM tends to outperform Mamba at larger scales (206M). This is an important advantage of xLSTM, as LRAM agents can strongly benefit from more data and consequently larger models. Note, that Mamba has a significantly higher number of parameters than competitors. For the zero-shot evaluation performances on the 37 hold-out tasks, we refer to Figure 14 in Appendix D.2.
Performance per domain. In Figure 3, we report the normalized scores for the 206M models attained on all six domains. For Meta-World, DMControl, Mimicgen, Composuite, and Procgen we use data-normalized scores, as suggested by (Levine et al., 2020). For Atari, we report human-normalized scores. We observe that the xLSTM backbone outperforms competitors on three of the six domains, while they perform similarly on the remaining domains.
@@ 4.3 Analyses & Ablations
Embedding space analysis. In Figure 5, we analyze the representations learned by our model. We sample 32 sub-trajectories from every task, extract the sequence representation at the last layer, cluster them using UMAP (McInnes et al., 2018), and color every point by its domain (see Appendix F for more details). We find that tasks from the same domain cluster together. Furthermore, xLSTM exhibits a more refined domain separation compared to DT, which may further contribute to the better down-stream performance.
@@
Figure 6: Latency comparison on A100. We report latency for varying context lengths (in timesteps) with batch sizes (a) $B=1$ and (b) $B=16$. In (c), we show the memory consumption in % of GPU memory with $B=1$. We compare DT to xLSTM with the same number of layer blocks and parameters on Atari Freeway. Missing bars for DT indicate out-of-memory (OOM).
Removing Actions & Effect of Context Length. We found that removing actions from the context results in better performance across backbones. While context lengths beyond 1 hurt performance on Meta-World and DMControl and when training with actions, the reverse is true when training without actions (see Figures 22, 23, 25). This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). While removing actions improves performance on Meta-World/DMControl, it does not affect performance on discrete control environments. For Meta-World/DMControl, we observed that the models become overly confident, which is problematic if poor initial actions are produced. This is because many robotics environments exhibit smoothly changing actions and by observing previous actions the agent learns shortcuts. A similar issue has been observed by Wen et al. (2020) and termed the copycat problem. Removing actions from the input prevents the agent from using shortcuts and alleviates the copycat problem. Importantly, the evaluation performance improves across domains as the sequence length increases, which indicates that the history helps to predict the next action (e.g., by observing mistakes made in the past, see Figures 24, 26).
Return-conditioning vs. Behavior Cloning. Across our experiments, we utilized a sequence representation that includes return-to-go tokens, as commonly used in DTs (Chen et al., 2021; Lee et al., 2022). However, many recent works focus on behavior cloning without return conditioning (Reed et al., 2022; Brohan et al., 2023a). Therefore, we study the effect of excluding the RTG/reward tokens from the sequence at the 206M parameter scale, to validate that our findings transfer to the behavior cloning setting. Indeed, we find that the same trends hold (see Figures 27 and 28).
@@ 4.4 Inference Time Comparison
Finally, we empirically examine the difference between xLSTM-based and Transformer-based agents at inference time. Similar to De et al. (2024), we report both latency and throughput. We focus our analysis on latency, as it is the more important dimension for real-time applications.
Setup. We conduct all inference time tests on A100s with 40GB of RAM using 206M models. For the Transformer, we use KV-caching and FlashAttention (Dao, 2023) as supported by PyTorch (Paszke et al., 2019). For xLSTM, we use recurrent-style inference using custom kernels to accelerate computations (see Figure 21 for the impact of kernel acceleration). For both backbones, we use torch.compile. The Transformer with KV-caching has a linear time complexity per step and quadratic in the sequence length. In contrast, the xLSTM has a constant time complexity per step and linear in the sequence length. Therefore, we expect speed-ups especially for longer sequences and larger batch sizes, as observed by De et al. (2024). To ensure a fair comparison, we compare DT and xLSTM with the same number of layer blocks and increase the hidden size of xLSTM to match the number of parameters of DT (see Appendix E.5 for evaluation performance of these models). We provide further details on our inference time tests in Appendix D.5.
@@
Latency. Similar to De et al. (2024), we measure latency by the average time (in seconds) taken to perform a single inference step with a fixed batch size $B$ (lower is better). In Figure, 6, we report the latencies for varying context lengths, $C\in[50,25600]$ and two batch sizes $B\in\{1,16\}$. Note that $C$ is in time steps and every time step contains 3 tokens (state, reward-to-go, reward). Hence, the effective sequence length for the largest $C$ is 76800. As expected, we find that the recurrent backbone attains lower inference latencies than the Transformer one. As the sequence length increases, DT runs out of memory due to the increasing size of the KV cache (see Figure 6c). In contrast, the inference speeds for xLSTM are independent of the context length, and therefore enable significantly longer context lengths. This property is particularly interesting for in-context RL, which requires keeping multiple episodes in the context (Laskin et al., 2022). Nevertheless, our experiments highlight that the materialization of the complexity advantage depends on the device, model size, batch size and the context length, which is similar to findings by De et al. (2024).
Figure 7: Throughput comparison on A100 for varying batch sizes with $C=1600$ timesteps on the Atari Freeway environment. Missing bars for DT indicate OOM.
Throughput. Throughput is measured by the total amount of inference steps performed per second for a model with a fixed context length. In Figure, 7, we report the throughputs for varying batch sizes, $B\in[1,128]$ for a fixed context length of $C=1600$. Here, the batch size can be interpreted as the number of parallel environments the agent interacts with. As expected, we find that xLSTM attains considerably higher throughputs than the DT. The benefit of xLSTM increases with larger batch sizes. While the DT with quadratic complexity in the sequence length goes OOM for batch sizes above 64, the xLSTM with linear complexity can easily handle larger batch sizes. In both experiments, the recurrent xLSTM performs favorably over the Transformer backbone.
@@ 5 Conclusion
In this work, we study the aptitude of modern recurrent architectures as alternatives to Transformers for building LAMs. We found that our LRAM with an xLSTM or Mamba at its core compare favorably to the Transformer in terms of evaluation performance across model scales (see Section 4.2). Moreover, we demonstrated that xLSTM-based LRAMs exhibit higher inference speeds, especially at large context sizes (see Section 4.4). Thus, the empirical evidence suggests, that recurrent backbones can be attractive alternatives for LAMs. Notably, the linear-time inference complexity of xLSTM may enable applications that require long context lengths (e.g., ICL), and facilitate the application of large-scale agents for real-time applications, such as robotics.
@@
Modern recurrent architectures and Transformers come with different pros and cons. xLSTM and Mamba, on the one hand, exhibit a complexity advantage over Transformers. Their linear complexity ensures that the computational requirements increase slower with the sequence length, which enables more efficient inference and is particularly relevant for edge-applications. While we conduct our inference time comparisons on a high-end data-center GPU, applications on edge-devices may have to deal with less powerful accelerators. Importantly, we found that LAMs strongly benefit from longer sequences (see Section 4.3). Transformers, on the other hand, are effective for applications that require exact recall of tokens in a sequence, which can be important for decision-making (Ni et al., 2024). Finally, xLSTM in particular enables state-tracking via sLSTM blocks, which Transformers and Mamba cannot perform (Merrill et al., 2024). State tracking can be important for logic tasks or dealing with partial observability and may be a useful tool for practitioners. Given these differences, different backbones should be considered depending on the task at hand.
@@
Limitations. The primary target application of LAMs is robotics. While the majority of our experiments involve robotic simulations, we do not yet provide experiments for real robots. We do, however, believe that our findings translate to real-world scenarios and aim to provide further evidence in future work. Moreover, our fine-tuning experiments are limited to offline RL. We envision that an agent pre-trained on large-scale datasets can be successfully fine-tuned via online RL to explore new strategies that do not appear in the training data. Modern recurrent architectures offer both parallel and recurrent training mode, which might be the key to success for such applications. While we provide evidence for improved ICL abilities of LRAM, we only consider a grid-world setting. We aim to further investigate the ICL abilities of LRAM in more complex environments.
@@
While we conduct all our experiments in simulated environments, the primary target application of our method is robotics. We believe that our work can positively impact applications in the near future, which require efficient inference, on-device processing, or have real-time constraints. However, robotics applications in the real world are not without risks. In particular, in areas where humans are involved, such as factory settings, special care is required. LAMs are trained via next-action prediction similar to LLMs. Consequently, LAMs may also suffer from hallucinations in unknown scenarios. We therefore strongly discourage users from blindly following the predictions made by real-world LAMs without appropriate safeguards regarding safety and robustness. It is essential to ensure responsible deployment of such future technologies, and we believe that more research on the robustness of LAMs is necessary.
@@ Acknowledgements
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE12]. We thank the projects INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
@@ Appendix A Reproducibility Statement
We make the code-base used for our experiments publicly available, and release the datasets we generated. Both are available at: https://github.com/ml-jku/LRAM. As part of this submission, we also include the source code in the supplementary material. We describe the environments we use for our experiments and provide dataset statistics in Appendix B. Furthermore, in Appendix C, we provide implementation details for all methods and a list of hyperparameters used for our experiments. In Appendix D, we present additional figures that accompany our results in the main text (e.g., all model sizes). Finally, in Appendices E and F, we provide further details on the conducted ablation studies and the embedding space analysis, respectively.
@@ B.1 General
We compile a large-scale dataset comprising 432 tasks from six domains, 3.4M trajectories, and 894M transitions in total (see Table 1). To enable fast and targeted data-loading, every trajectory is stored in a separate hdf5 file. We trade off some data-loading speed for disk space efficiency, by compressing trajectories that contain image-based observations.
@@ B.2 Atari
The Arcade Learning Environment (ALE) (Bellemare et al., 2013) is the standard benchmark for evaluating RL agents and consists of 57 Atari games. Input observations in Atari are RGB images, but as is standard practice we gray-scale and crop frames ($|\mathcal{S}|=1\times 64\times 64$). There are 18 discrete action across all 57 Atari games ($|\mathcal{A}|=18$), but individual games may use only use a subset of these actions. Furthermore, we adopt the standard Atari recipe as used in prior works, including a frame skip of 4, maximum number of no-ops of 30, resetting on life loss, and reward clipping to $[-1,1]$ (Mnih et al., 2015; Hessel et al., 2017).
Tasks. Similar to Lee et al. (2022), we assign 41 games to the training set, and 5 additional tasks to the hold-out set. The 41 training tasks include:
@@
Dataset. For Atari, we leverage the DQN-Replay dataset released by Agarwal et al. (2020). The dataset contains the trajectories seen over the entire training of the DQN agent (50M frames), We extract a subset of the last 5M transitions for every task, amounting to 205M transitions in total for the 41 training tasks. The number of episodes, the episodes lengths and total achieved rewards vary across tasks, as shown in Table 2.

### v3

@@ 4.2 Scaling comparison
To conduct our main comparisons, we train our four backbone variants on the full training task mixture of 432 tasks. For each architecture backbone, we report performance scores for four model sizes: 16M, 48M, 108M, and 206M parameters. We train all models for 200K updates with a batch size of 128 and a context length of 50 timesteps. All domains are represented with approximately equal proportion, resulting in 33K updates per domain. Additional implementation details and hyperparameters for every backbone variant and model size are available in Appendix C.
@@
Evaluation performance. During training, we evaluate our agents after every 50K step in all 432 training environments. In Figure 2b, we report the resulting normalized performances averaged across all six domains. The recurrent backbones outperform the Transformer one across model sizes. While xLSTM and Mamba perform similarly at smaller scales, xLSTM tends to outperform Mamba at larger scales (206M). This is an important advantage of xLSTM, as LRAM agents can strongly benefit from more data and consequently larger models. Note that Mamba has a significantly higher number of parameters than competitors. For the zero-shot evaluation performances on the 37 hold-out tasks, we refer to Figure 14 in Appendix D.2.
Performance per domain. In Figure 3, we report the normalized scores for the 206M models attained on all six domains. For Meta-World, DMControl, Mimicgen, Composuite, and Procgen, we use data-normalized scores, as suggested by (Levine et al., 2020). For Atari, we report human-normalized scores. We observe that xLSTM outperforms competitors on three of the six domains, while they perform similarly on the remaining domains.
@@ 4.3 Analyses & Ablations
Embedding space analysis. In Figure 5, we analyze the representations learned by our model. We sample 32 sub-trajectories from every task, extract the sequence representation at the last layer, cluster them using UMAP (McInnes et al., 2018), and color every point by its domain (see Appendix F for more details). We find that tasks from the same domain cluster together. Furthermore, xLSTM exhibits a more refined domain separation compared to DT, which may further contribute to the better downstream performance. See Appendix F for a more detailed discussion on the embedding space analysis and a comparison to Mamba.
@@
Figure 6: Latency comparison on A100. We report latency for varying context lengths (in timesteps) with batch sizes (a) $B=1$ and (b) $B=16$. In (c), we show the memory consumption in % of GPU memory with $B=1$. We compare DT to xLSTM and Mamba with the same number of layer blocks and parameters on Atari Freeway. Missing bars for DT indicate out-of-memory (OOM).
Removing Actions & Effect of Context Length. We found that removing actions from the context results in better performance across backbones. While context lengths beyond 1 hurt performance on Meta-World and DMControl, and when training with actions, the reverse is true when training without actions (see Figures 23, 24, 26). This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). While removing actions improves performance on Meta-World/DMControl, it does not affect performance on discrete control environments. For Meta-World/DMControl, we observed that the models become overly confident, which is problematic if poor initial actions are produced. This is because many robotics environments exhibit smoothly changing actions, and by observing previous actions, the agent can learn shortcuts. A similar issue has been observed by Wen et al. (2020) and termed the copycat problem. Removing actions from the input prevents the agent from using shortcuts and, therefore, alleviates the copycat problem. Importantly, the evaluation performance improves across domains as the sequence length increases, which indicates that the history helps to predict the next action (e.g., by observing mistakes made in the past, see Figures 25, 27).
Return-conditioning vs. Behavior Cloning. Across our experiments, we utilized a sequence representation that includes return-to-go tokens, as commonly used in DTs (Chen et al., 2021; Lee et al., 2022). However, many recent works focus on behavior cloning without return conditioning (Reed et al., 2022; Brohan et al., 2023a). Therefore, we study the effect of excluding the RTG/reward tokens from the sequence at the 206M parameter scale, to validate that our findings transfer to the behavior cloning setting. Indeed, we find that the same trends hold (see Figures 28 and 29).
@@ 4.4 Inference Time Comparison
Finally, we empirically examine the difference between recurrent and Transformer-based agents at inference time. Similar to De et al. (2024), we report both latency and throughput. We focus our analysis on latency, as it is the more important dimension for real-time applications.
Setup. We conduct all inference time tests on A100s with 40GB of RAM using 206M models. For the Transformer, we use KV-caching and FlashAttention (Dao, 2023) as supported by PyTorch (Paszke et al., 2019). For xLSTM, we use recurrent-style inference using custom kernels to accelerate computations (see Figure 21 for the impact of kernel acceleration). For Mamba, we make use of the kernels introduced by Gu & Dao (2023). For DT and xLSTM, we use torch.compile, but not for Mamba because we found the kernels to be incompatible with compilation. The Transformer with KV-caching has a linear time complexity per step and quadratic in the sequence length. In contrast, the xLSTM and Mamba have a constant time complexity per step and are linear in the sequence length. Therefore, we expect speed-ups especially for longer sequences and larger batch sizes, as observed by De et al. (2024). To ensure a fair comparison, we compare all backbones with the same number of layer blocks and increase the hidden size of xLSTM and Mamba to match the number of parameters of DT (see Appendix E.5 for evaluation performance of these models). We provide further details on our inference time tests in Appendix D.5.
@@
Latency. Similar to De et al. (2024), we measure latency by the average time (in seconds) taken to perform a single inference step with a fixed batch size $B$ (lower is better). In Figure, 6, we report the latencies for varying context lengths, $C\in[50,25600]$ and two batch sizes $B\in\{1,16\}$. Note that $C$ is in time steps, and every time step contains 3 tokens (state, reward-to-go, reward). Hence, the effective sequence length for the largest $C$ is 76800. As expected, we find that the recurrent backbones attain lower inference latencies than the Transformer one, especially for longer sequences and with a larger batch size. For $B=1$, we find that Mamba is slower than the Transformer and xLSTM, which we believe is because of the incompatibility with torch.compile. Note that we expect the gap to xLSTM to be closed with compatible kernels. As the sequence length increases, DT runs out of memory due to the increasing size of the KV cache (see Figure 6c). In contrast, the inference speeds for Mamba/xLSTM are independent of the context length and therefore, enable significantly longer context lengths. This property is particularly interesting for in-context RL, which requires keeping multiple episodes in the context (Laskin et al., 2022). Nevertheless, our experiments highlight that the materialization of the complexity advantage depends on the device, model size, batch size, and the context length, which is similar to findings by De et al. (2024).
Figure 7: Throughput comparison on A100 for varying batch sizes with $C=1600$ timesteps on the Atari Freeway environment. We compare DT, xLSTM with 4 and 16 heads, and Mamba. Missing bars for DT indicate OOM.
Throughput. Throughput is measured by the total number of inference steps performed per second for a model with a fixed context length. In Figure 7, we report the throughputs for varying batch sizes, $B\in[1,128]$ for a fixed context length of $C=1600$. Here, the batch size can be interpreted as the number of parallel environments the agent interacts with. For xLSTM, we report numbers for two variants with 4 and 16 heads, respectively. We found that decreasing the head dimension (more heads, same total hidden dim) is important for xLSTM to enable high throughput. This is because a higher head dimension incurs more FLOPS (see Figure 22 in Appendix D.5.4 for an ablation on the impact of the head dimension). As expected, we find that both Mamba and xLSTM attain considerably higher throughputs than the DT. These benefits increase with larger batch sizes. While the DT with quadratic complexity in the sequence length goes OOM for batch sizes above 64, the recurrent backbones with linear complexity can easily handle larger batch sizes. This throughput advantage may be particularly relevant for online fine-tuning of agents in many parallel environments.
@@ 5 Conclusion
In this work, we study the aptitude of modern recurrent architectures as alternatives to Transformers for building LAMs. We found that our LRAM with an xLSTM or Mamba at its core compares favorably to the Transformer in terms of evaluation performance across model scales ranging from 16M to 206M parameters (see Section 4.2). Moreover, we demonstrated that LRAM exhibits higher inference speeds, especially at large context sizes (see Section 4.4). Thus, the empirical evidence suggests that recurrent backbones can be attractive alternatives for LAMs. Notably, the linear-time inference complexity of xLSTM and Mamba may enable applications that require long context lengths (e.g., ICL) and facilitate the application of large-scale agents for real-time applications, such as robotics.
@@
Modern recurrent architectures and Transformers come with different advantages and disadvantages. xLSTM and Mamba, on the one hand, exhibit a fundamental complexity advantage over Transformers. Their linear complexity ensures that the computational requirements increase more slowly with the sequence length, which enables more efficient inference and is particularly relevant for edge applications. While we conduct our inference time comparisons on a high-end data center GPU, applications on edge devices may have to deal with less powerful accelerators. Importantly, we found that LAMs strongly benefit from longer sequences (see Section 4.3). Their ability to efficiently handle long sequences can be beneficial for applications in real-world environments, which often exhibit long-term dependencies. Similarly, longer context can be relevant for ICL applications, which benefit from keeping multiple episodes (such as demonstrations or previous trials) in the context. Transformers, on the other hand, are effective for applications that require exact recall of tokens (such as particular locations in a grid, signs in an image) in a sequence, which can be important for decision-making (Ni et al., 2024). Finally, xLSTM in particular enables state-tracking via sLSTM blocks, which Transformers and Mamba cannot perform (Merrill et al., 2024). State tracking can be important for logic tasks or dealing with partial observability and may be a useful tool for practitioners. Given these differences, different backbones should be considered depending on the task at hand.
@@
Limitations & Future Work. The primary target application of LAMs is robotics. While the majority of our experiments involve robotic simulations, we do not yet provide experiments for real robots. We do, however, believe that our findings translate to real-world scenarios and aim to provide further evidence in future work. Moreover, our fine-tuning experiments are limited to offline RL. We envision that an agent pre-trained on large-scale datasets can be successfully fine-tuned via online RL to explore new strategies that do not appear in the training data. Modern recurrent architectures offer both parallel and recurrent training modes, which might be the key to success for such applications. While we provide evidence for improved ICL abilities of LRAM, we only consider a grid-world setting. We aim to further investigate the ICL abilities of LRAM in more complex environments.
@@
While we conduct all our experiments in simulated environments, the primary target application of our method is robotics. We believe that our work can positively impact applications in the near future that require efficient inference, on-device processing, or have real-time constraints. However, robotics applications in the real world are not without risks. In particular, in areas where humans are involved, such as factory settings, special care is required. LAMs are trained via next-action prediction similar to LLMs. Consequently, LAMs may also suffer from hallucinations in unknown scenarios. We therefore strongly discourage users from blindly following the predictions made by real-world LAMs without appropriate precautions regarding safety and robustness. It is essential to ensure the responsible deployment of such future technologies, and we believe that more research on the robustness of LAMs is necessary.
@@ Acknowledgements
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG.
@@ Appendix A Reproducibility Statement
We make the code base used for our experiments publicly available and release the datasets we generated. Both are available at: https://github.com/ml-jku/LRAM. We describe the environments we use for our experiments and provide dataset statistics in Appendix B. Furthermore, in Appendix C, we provide implementation details for all methods and a list of hyperparameters used for our experiments. In Appendix D, we present additional figures that accompany our results in the main text (e.g., all model sizes). Finally, in Appendices E and F, we provide further details on the conducted ablation studies and the embedding space analysis, respectively.
@@ B.1 General
We compile a large-scale dataset comprising 432 tasks from six domains, 3.4M trajectories, and 894M transitions in total (see Table 1). A key motivation behind our dataset compilation is the scarcity of suitable datasets that span many simulated tasks. To address this and to enable a robust comparison of different sequence model architectures, we aimed to assemble a collection of datasets that span as many tasks as possible. In particular, we focused on trajectories in simulated environments rather than real-world trajectories (Embodiment Collaboration et al., 2024), to enable faster iteration cycles. To facilitate usability for future works, we consider standard benchmarks that are widely adopted by the community (e.g., Atari, Meta-World).
We release our data pipeline and generated dataset, and hope that they can serve as a solid basis for future research on multi-task agents. To enable fast and targeted data-loading, every trajectory is stored in a separate hdf5 file. We trade off some data-loading speed for disk space efficiency by compressing trajectories that contain image-based observations.
@@ B.2 Atari
The Arcade Learning Environment (ALE) (Bellemare et al., 2013) is the standard benchmark for evaluating RL agents and consists of 57 Atari games. Input observations in Atari are RGB images, but as is standard practice, we gray-scale and crop frames ($|\mathcal{S}|=1\times 64\times 64$). There are 18 discrete actions across all 57 Atari games ($|\mathcal{A}|=18$), but individual games may use only a subset of these actions. Furthermore, we adopt the standard Atari recipe as used in prior works, including a frame skip of 4, maximum number of no-ops of 30, resetting on life loss, and reward clipping to $[-1,1]$ (Mnih et al., 2015; Hessel et al., 2017).
Tasks. Similar to Lee et al. (2022), we assign 41 games to the training set and 5 additional tasks to the hold-out set. The 41 training tasks include:
@@
Dataset. For Atari, we leverage the DQN-Replay dataset released by Agarwal et al. (2020). The dataset contains the trajectories seen over the entire training of the DQN agent (50M frames). We extract a subset of the last 5M transitions for every task, amounting to 205M transitions in total for the 41 training tasks. The number of episodes, the episode lengths, and total achieved rewards vary across tasks, as shown in Table 2.